OpenStudy (anonymous):

find the real-number solution... p^3+4p^2-9p=0

OpenStudy (lyrae):

\[p(p^2+4p-9) = 0\]p=0 \[p^2+4p-9 =0\]\[-4/2 \pm \sqrt{(4/2)^{2} - (-9)}\]\[-2 \pm \sqrt{13} \]p = -2 - sqrt(13) p=sqrt(13) - 2